OpenStudy (anonymous):

(SOLVE! BEST ANSWER GETS MEDAL! SHOW WORK!) [(y + 1)/4] - [(y - 1)/6] = y + 5

OpenStudy (anonymous):

Simplify brackets y+14−y−16=y+5 2 Multiply both sides by 12 (the LCM of 4,6) 3(y+1)−2(y−1)=12y+60 3 Expand 3y+3−2y+2=12y+60 4 Simplify 3y+3−2y+2 to 3y−2y+5 3y−2y+5=12y+60 5 Simplify 3y−2y+5 to y+5 y+5=12y+60 6 Subtract y from both sides 5=12y+60−y 7 Simplify 12y+60−y to 11y+60 5=11y+60 8 Subtract 60 from both sides 5−60=11y 9 Simplify 5−60 to −55 −55=11y 10 Divide both sides by 11 −5511=y 11 Simplify 5511 to 5 −5=y 12 Switch sides y=−5

OpenStudy (anonymous):

DANG IT!! Sorry

OpenStudy (anonymous):

Lol what are all the question marks? xD

OpenStudy (anonymous):

Idk

OpenStudy (anonymous):

Simplify brackets y+14−y−16=y+5 2 Multiply both sides by 12 (the LCM of 4,6) 3(y+1)−2(y−1)=12y+60 3 Expand 3y+3−2y+2=12y+60 4 Simplify 3y+3−2y+2 to 3y−2y+5 3y−2y+5=12y+60 5 Simplify 3y−2y+5 to y+5 y+5=12y+60 6 Subtract y from both sides 5=12y+60−y 7 Simplify 12y+60−y to 11y+60 5=11y+60 8 Subtract 60 from both sides 5−60=11y 9 Simplify 5−60 to −55 −55=11y 10 Divide both sides by 11 −5511=y 11 Simplify 5511 to 5 −5=y 12 Switch sides y=−5 There we go!!

OpenStudy (anonymous):

I hope that helps you out

OpenStudy (anonymous):

The numbers are the steps btw

OpenStudy (anonymous):

Okay thank you! :)

OpenStudy (anonymous):

welcome